# Pseudo-real quantum fields

### v1

@@ Appendix A Calculating the $\hat{\eta}$ operator
We note that the expansion of the $\eta_{P}$ transformation in Eq. (3.29) has the same terms as the expansion of $\phi$ in Eq. (3.10). Hence, we make an Ansatz that the operator $\hat{\eta}_{P}$ should be of the form of the time-evolution operator, but with differing coefficients and expressed in terms of the full, interacting field $\hat{\phi}$. We define an operator $\hat{U}$
$\hat{U}\quantity(\alpha,t,t_{0})\coloneqq T\quantity[e^{\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\text{\>\>\>and\>\>\>}\hat{U}^{-1}\quantity(\alpha,t,t_{0})=\bar{T}\quantity[e^{-\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (A.1)
for some constant $\alpha\in\mathbb{C}$. Here, $T$ is the time-ordering operator and $\bar{T}$ is the anti-time ordering operator.
@@
[…]
We note that the commutator $[\hat{\phi}_{x},\hat{\phi}_{y}]$ corresponds to that of the interacting field $\hat{\phi}$ and not to that of the free Klein-Gordon field $\hat{\phi}_{0}$. However, we can expand the interacting field as a series in the coupling constant $g$. For our purposes, we only need to perform this expansion at first order in $g$:
@@
We calculate the commutators for the expansion in Eq. (A.2) at order $g$:
@@
For the third term in Eq. (A.2), we only need the commutator at 00-th order in g:
$[\hat{\phi}^{3}_{z},[\hat{\phi}^{3}_{y},\hat{\phi}_{x}]]=-18G_{xy}G_{yz}\hat{\phi}_{y}\hat{\phi}^{2}_{z}+18iG_{xy}G_{yz}G_{yz}\hat{\phi}_{z}\;.$ (A.8)
@@
Finally, taking $\alpha=-\frac{1}{3}$, we get the operator $\hat{\eta}_{g}(t)=\hat{U}^{-1}\quantity(-\frac{1}{3},t,t_{0})$ in Eq. (3.6)
@@
with the last term appearing due to quantum loop corrections. It is proportional to $\hbar$ and vanishes in the classical limit $\hbar\mapsto 0$.
@@ Appendix B Proving the Hermiticity of $\hat{\eta}$
The pseudo-reality condition gives us an expansion of $\hat{\phi}^{\dagger}$ in terms of $\hat{\phi}$ [see Eq. (3.30)]. At order $g$, we have
@@
Hence, the interaction term, commuting all $\hat{\phi}_{x}$ to the left of $\hat{\phi}_{y}$, is
@@
$\hat{\mathcal{P}}^{\dagger}(t)=e^{\frac{\pi}{2}\int d^{3}z\>\hat{\phi}^{\dagger}(t,\vec{z})\hat{\pi}^{\dagger}(t,-\vec{z})+\hat{\phi}^{\dagger}(t,\vec{z})\hat{\pi}^{\dagger}(t,\vec{z})}\;,$ (B.9)
@@
Using the expansion of $\hat{\phi}^{\dagger}$ in terms of the field $\hat{\phi}$, we can show that $\hat{\pi}^{\dagger}$ satisfies an analogous pseudo-reality condition:
@@
Finally, we can use the above results to show that $\hat{\eta}_{P}=\hat{\eta}_{g}\hat{\mathcal{P}}$ is Hermitian:
$\hat{\eta}_{P}^{\dagger}=\hat{\mathcal{P}}^{\dagger}\hat{\eta}_{g}^{\dagger}=\hat{\eta}_{P}\hat{\mathcal{P}}\hat{\eta}_{P}^{-1}\hat{\eta}_{g}=\hat{\eta}_{P}\hat{\mathcal{P}}\hat{\mathcal{P}}^{-1}\hat{\eta}_{g}^{-1}\hat{\eta}_{g}=\hat{\eta}_{P}\;.$ (B.12)

### v2

@@ Appendix A Calculating the $\hat{\eta}$ operator
We note that the expansion of the $\eta_{P}$ transformation in Eq. (47) has the same terms as the expansion of $\phi$ in Eq. (26). Hence, we make an Ansatz that the operator $\hat{\eta}_{P}$ should be of the form of the time-evolution operator, but with differing coefficients and expressed in terms of the full, interacting field $\hat{\phi}$. We define an operator $\hat{U}$
$\hat{U}\quantity(\alpha,t,t_{0})\coloneqq T\quantity[e^{\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (A.90)
and
$\hat{U}^{-1}\quantity(\alpha,t,t_{0})=\bar{T}\quantity[e^{-\alpha g\int^{t}_{t_{0}}{\rm d}y^{0}\int{\rm d}^{3}\vec{y}\>\hat{\phi}^{3}(y^{0},\vec{y})}]\>,$ (A.91)
for some constant $\alpha\in\mathbb{C}$. Here, $T$ is the time-ordering operator and $\bar{T}$ is the anti-time-ordering operator.
@@
where we define shorthand notations for the integral $\int_{t_{0},\>\vec{y}}^{x^{0}}\>\coloneqq\int_{t_{0}}^{x^{0}}{\rm d}y^{0}\>\int{\rm d}^{3}\vec{y}$ and the field $\hat{\phi}_{x}\coloneqq\hat{\phi}(x^{0},\vec{x})$.
We note that the commutator $\hat{\phi}_{x},\hat{\phi}_{y}]$ corresponds to that of the interacting field $\hat{\phi}$ and not to that of the free Klein-Gordon field $\hat{\phi}_{0}$. However, we can expand the interacting field as a series in the coupling constant $g$. For our purposes, we only need to perform this expansion at first order in $g$:
@@
We calculate the commutators for the expansion in Eq. (A.92) at order $g$ to obtain
@@
For the third term in Eq. (A.92), we only need the commutator at 0th order in g:
$\hat{\phi}^{3}_{z},[\hat{\phi}^{3}_{y},\hat{\phi}_{x}]]=-18G_{xy}G_{yz}\hat{\phi}_{y}\hat{\phi}^{2}_{z}+18iG_{xy}G_{yz}G_{yz}\hat{\phi}_{z}\;.$ (A.98)
@@
Finally, taking $\alpha=-\frac{1}{3}$, we get the operator $\hat{\eta}_{g}(t)=\hat{U}^{-1}\quantity(-\frac{1}{3},t,t_{0})$ in Eq. (III.6), giving
@@
with the last term appearing due to quantum loop corrections. It is proportional to $\hbar$ and vanishes in the classical limit $\hbar\mapsto 0$ (see Section III.6).
@@ Appendix B Proving the Hermiticity of $\hat{\eta}$
The pseudo-reality condition gives us an expansion of $\hat{\phi}^{\dagger}$ in terms of $\hat{\phi}$ [see Eq. (48)]. At order $g$, we have
@@
Hence, after commuting all $\hat{\phi}_{x}$ to the left of $\hat{\phi}_{y}$, the interaction term is
@@
$\hat{\mathcal{P}}^{\dagger}(t)=e^{\frac{\pi}{2}\int{\rm d}^{3}\vec{z}\>\hat{\phi}^{\dagger}(t,\vec{z})\hat{\pi}^{\dagger}(t,-\vec{z})+\hat{\phi}^{\dagger}(t,\vec{z})\hat{\pi}^{\dagger}(t,\vec{z})}\;,$ (B.109)
@@
Using the expansion of $\hat{\phi}^{\dagger}$ in terms of $\hat{\phi}$, we can show that $\hat{\pi}^{\dagger}$ satisfies an analogous pseudo-reality condition, with
@@
Finally, we can use the above results to show that $\hat{\eta}_{P}=\hat{\eta}_{g}\hat{\mathcal{P}}$ is Hermitian, viz.,
$\hat{\eta}_{P}^{\dagger}=\hat{\mathcal{P}}^{\dagger}\hat{\eta}_{g}^{\dagger}=\hat{\eta}_{P}\hat{\mathcal{P}}\hat{\eta}_{P}^{-1}\hat{\eta}_{g}=\hat{\eta}_{P}\hat{\mathcal{P}}\hat{\mathcal{P}}^{-1}\hat{\eta}_{g}^{-1}\hat{\eta}_{g}=\hat{\eta}_{P}\;,$ (B.112)
as required.